what is the coefficent A,B,C, in the quadratic equation
2x^2-5x=1?

Answer by RAY100(1637)

basic form is Ax^2+Bx+C=0 for using quadratic eqn to solve for x
your eqn is 2x^2-5x =1
subt 1 both sides
(2)x^2 +(-5)x +(-1) =0
therefore A=(2), B=(-5), C=(-1)
standard form of quadratic eqn is Ax^2+Bx=C
in which case a=(2), B=(-5), but C=(+1)
check which problem you have